Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 564
Entry time: 11 Nov 2024, 09:46
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 11.11.2024 |
|
|
General Announcement:
- Last day of the FD taking place and switching back to 16.3GeV today.
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- Nothing to report.
VAC:
- Nothing to report.
XPD:
- FD studies went fine during the weekend; Attoseconds went fine; XFELO progressed.
- PES can be used for the remaining delivery weeks for 2CPP but not for polarization.
DET:
- Nothing major to report.
ITDM:
- Nothing special to report.
XRO:
- XFELO: Alignment problems as expected; first time using 4 undulators from inside the cavity; pointing/trajectory inside had changed and to be re-done on Sunday.
- In general, all studies brought some results even if not majors.
LAS:
- Not present.
UNSYS:
- Nothing to report.
- Polarized attoseconds need more time in FD weeks.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
MID:
- Nothing to report.
HED:
- Setting up for two programs this week; all fine so far.
SCS:
- Nothing to report.
SQS:
- Take beam this week for IHR; finishing preparation.
SXP:
- Not present. |